A self-contained underwater breathing apparatus uses canisters containing potassium superoxide. The superoxide consumes the exhaled by a person and replaces it with oxygen.
4 KO2(s)+2 CO2(g)--> 2 K2CO3(s)+3 O2(g)
What mass of KO2 , in grams, is required to react with 8.40 L of CO2 at 23.0°C and 764 mm Hg?

Mass =____ g
1
Expert's answer
2020-05-03T15:14:02-0400

Solution:

According to ideal gas equation: PV = nRT.

P = 764 mmHg = 1.00526 atm.

V = 8.40 L;

R = 0.08206 L atm K-1 mol-1;

T = 23.0°C = 296.15 K.

Then,

n = PV / RT

n(CO2) = (1.00526 atm × 8.40 L) / (0.08206 L atm K-1 mol-1 × 296.15 K) = 0.3475 mol

n(CO2) = 0.3475 mol.


The balanced chemical equation:

4KO2(s) + 2CO2(g) --> 2K2CO3(s) + 3O2(g) 

According to chemical equation: n(KO2)/4 = n(CO2)/2;

n(KO2) = 4×n(CO2)/2 = 2×n(CO2).

n(KO2) = 2 × (0.3475 mol) = 0.695 mol.


Mass of KO2 = Moles of KO2 × Molar mass of KO2

The molar mass of KO2 is 71.1 g/mol.

Then,

Mass of KO2 = (0.6948 mol) × (71.1 g/mol) = 49.4145 g = 49.41 g

m(KO2) = 49.41 g.


Answer: Mass of KO2 is 49.41 g.